OpenStudy (metalslayer):

Use the first and last points to find algebraically the particular equation of the natural logarithmic function that fits the point

OpenStudy (metalslayer):

Table: x= 3.6, 14.4, 57.6, 230.4, 921.6 y= 1, 2, 3, 4, 5

OpenStudy (metalslayer):

\[y=\log_{4} \frac{ 10 }{ 9 }x\]

OpenStudy (metalslayer):

That is the answer But how do I get the answer? Thanks!

myininaya (myininaya):

did you try to enter the last and first points into \[y=\log_a(bx)\] and solve for a and b?

OpenStudy (metalslayer):

But how do I do that. My teacher didn't teach logarithms last year in Algebra 2, that's why this concept is new to me

myininaya (myininaya):

you can plug in the points \[(3.6,1) \text{ and } (921.6,5)\] since it said plug and the first and last points

OpenStudy (snowsurf):

logarithmic is the inverse of a exponent.

myininaya (myininaya):

(x,y) replace x in my equation with the first number of the pair and replace y in my equation with the second number of the pair you will have two equations

OpenStudy (snowsurf):

@myininaya you will have to tell him what a logarithm is because he said he never had it in a previous course.

OpenStudy (metalslayer):

I understand what logarithm is, but I never performed any operations on them, that's why this is confusing for me

OpenStudy (metalslayer):

Will I need to create two equations for this problem?

OpenStudy (metalslayer):

\[1=\log_{a}(3.6x)\]

OpenStudy (metalslayer):

\[5=\log_{a}(921.6b) \]

OpenStudy (metalslayer):

like this? except the first equation: x should be b

OpenStudy (snowsurf):

Okay. So you have two equations. How would you simplify this? Any ideas? Use this relation \[ a^y=x\] \[y = \log_{a} x\]

OpenStudy (metalslayer):

I think by dividing the a's

OpenStudy (snowsurf):

You log equations should be setup like \[1 = \log_{a} (3.6x)\] \[5= \log_{a} (921.6x)\]

OpenStudy (metalslayer):

shouldn't it be 3.6b instead of 3.6x

OpenStudy (snowsurf):

Yes you right my bad. Since we don't know b.

OpenStudy (snowsurf):

Let's get of the logs \[1 = \log_{a} (3.6b)\] \[5 = \log_{a} (921.6b)\]

OpenStudy (metalslayer):

I understand the setup now, but what do I do next?

OpenStudy (metalslayer):

solve for a?

OpenStudy (snowsurf):

Yes. You can solve for either a or b.

OpenStudy (metalslayer):

Will logs cancel out? so it will be: a^1=3.6b and a^5= 921.6b

OpenStudy (snowsurf):

That is correct. Since you raise base a the logs are now gone. Now you need to solve for one of these variables

OpenStudy (metalslayer):

so it will be: a^-4 = (3.6/921.6)b

OpenStudy (metalslayer):

or do I need to flip the two equations

OpenStudy (snowsurf):

That will work but b should go away since your dividing two equations.

OpenStudy (snowsurf):

a^-4 = 3.6/921.6

OpenStudy (metalslayer):

And how to I get rid of the negative exponent?

OpenStudy (snowsurf):

Same as you would do with a non negative exponent. In this case you will take the negative root.

OpenStudy (metalslayer):

I get it, I just need to put 1/a^4

OpenStudy (snowsurf):

You can do that or \[a=\sqrt[-4]{3.6/921.6} \]

OpenStudy (metalslayer):

a^4 = 256 a=4 Thanks!

OpenStudy (snowsurf):

Now we can find b.

OpenStudy (metalslayer):

and after I plug in 4 for B in one of the equations.\[1=\log_{a} (3.6b)\]

OpenStudy (metalslayer):

\[1=\log_{a}(14.4) \]

OpenStudy (metalslayer):

then a = log 14.4

OpenStudy (metalslayer):

a=1.1584

OpenStudy (snowsurf):

Actually you can use your previous result to get b. Early you got \[a^1 = 3.6b\] Since you know a you can plug a in here and solve for b.

OpenStudy (metalslayer):

4=3.6b b= 1.1

OpenStudy (snowsurf):

Good. You got b now.

OpenStudy (metalslayer):

to get rid of the decimal, I can multiply the fraction's numerator and denominator by 10

OpenStudy (metalslayer):

and then reduce it to 10/9

OpenStudy (snowsurf):

That is correct. Good job.

OpenStudy (metalslayer):

\[y=\log_{4}\left(\begin{matrix}10 \\ 9\end{matrix}\right)x\]

OpenStudy (metalslayer):

(10/9x)